A book case is to have four shelves including the top. The height of the book case is to be 4.25 feet more than the width. Find the width and height of the book case if only 31 feet of lumber is available.
Answer by checkley77(12844)

2(4.25+x)+4x=31
8.5+2x+4x=31
2x+4x=31-8.5
6x=22.5
x=22.5/6
x=3.75 ft. is the width.
4.25+3.75=8 ft is the height.
Proof:
2(4.25+3.75)+4*3.75=31
2*8+15=31
16+15=31
31=31